# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

### v1

@@ 4.8 Endemic Equilibrium Analysis
Theorem 4.5 (Existence of Endemic Equilibrium)
The fractional-order Ebola system (2) admits a unique endemic equilibrium point $\mathcal{E}^{*}=(S^{*},V^{*},E^{*},I_{s}^{*},I_{a}^{*},H^{*},D^{*},R^{*})$ with all components strictly positive if and only if $\mathcal{R}_{0}>1$.
Proof.  At equilibrium, the time derivatives vanish. The force of infection is:
$\lambda^{*}=\frac{c\beta(I_{s}^{*}+\eta_{a}I_{a}^{*}+\eta_{d}D^{*})}{N^{*}}$
The equilibrium equations provide:
$0$ $=\Lambda-\lambda^{*}S^{*}+\omega V^{*}-q_{0}S^{*}$ (10)
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-q_{1}V^{*}$ (11)
$0$ $=\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]-q_{2}E^{*}$ (12)
$0$ $=p\sigma E^{*}-q_{3}I_{s}^{*}$ (13)
$0$ $=(1-p)\sigma E^{*}-q_{4}I_{a}^{*}$ (14)
$0$ $=h_{s}I_{s}^{*}-q_{5}H^{*}$ (15)
$0$ $=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}-q_{6}D^{*}$ (16)
$0$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}-\mu R^{*}$ (17)
Expressing infected states in terms of $E^{*}$ as:
$I_{s}^{*}$ $=\frac{p\sigma}{q_{3}}E^{*}\ and\quad I_{a}^{*}=\frac{(1-p)\sigma}{q_{4}}E^{*}$
$H^{*}$ $=\frac{p\sigma h_{s}}{q_{3}q_{5}}E^{*}and\quad D^{*}=\frac{p\sigma}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)E^{*}$
The force of infection becomes:
$\lambda^{*}=\frac{c\beta\sigma E^{*}}{N^{*}}\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
Define the composite parameter:
$M=c\beta\sigma\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
yielding $\lambda^{*}=ME^{*}/N^{*}$. From the exposed compartment equilibrium:
$E^{*}=\frac{\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]}{q_{2}},$
Uniting these relations gives the key equation:
$1=\frac{M}{q_{2}N^{*}}[S^{*}+(1-\varepsilon)V^{*}]$
Solving the susceptible-vaccinated subsystem:
$S^{*}$ $=\frac{\Lambda(q_{1}+(1-\varepsilon)\lambda^{*})}{\Delta(\lambda^{*})},\quad V^{*}=\frac{\Lambda v}{\Delta(\lambda^{*})}$
where $\Delta(\lambda^{*})=\left[(\omega+\mu+(1-\varepsilon)\lambda^{*})(\mu+v+\lambda^{*})\right]-\omega v$. This leads to $N^{*}=\Lambda/\mu$ hence the key equation becomes:
$1=\left[\frac{M}{q_{2}(\Lambda/\mu)}\right][S^{*}+(1-\varepsilon)V^{*}]$
At the disease-free state ($\lambda^{*}=0$), this expression equals $\mathcal{R}_{0}$. The monotonic decrease of $S^{*}+(1-\varepsilon)V^{*}]$ with $\lambda^{*}$ guarantees a unique positive solution $\lambda^{*}>0$ when $\mathcal{R}_{0}>1$, establishing the endemic equilibrium. $\square$
@@ 4.9 Local Stability Analysis of the Endemic Equilibrium
Consider system (2). When $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ is locally asymptotically stable.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}^{*}$ exhibits the block structure:
$J(\mathcal{E}^{*})=\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}$
with components:
$J_{11}$ $=\begin{pmatrix}-(\lambda^{*}+q_{0})&\omega\\ v&-\lambda^{*}+\varepsilon\lambda^{*}-q_{1})\end{pmatrix}$,
@@
$J_{22}$ $=\begin{pmatrix}-q_{2}&\beta\Psi\Phi&\beta\eta_{a}\Psi\Phi&0&\beta\eta_{d}\Psi\Phi&0\\ p\sigma&-q_{3}&0&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0&0\\ 0&h_{s}&0&-q_{5}&0&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}&0\\ 0&\gamma_{s}&\gamma_{a}&\gamma_{h}&0&-\mu\end{pmatrix}$
where $\Psi=S^{*}+(1-\varepsilon)V^{*}$, $\Phi=c/N^{*}$, and $\lambda^{*}$ denotes the equilibrium force of infection. The characteristic polynomial is an eighth-degree equation:
$P(\lambda)=\lambda^{8}+a_{7}\lambda^{7}+a_{6}\lambda^{6}+a_{5}\lambda^{5}+a_{4}\lambda^{4}+a_{3}\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0$
Consider the Metzler matrix $Q=-J(\mathcal{E}^{*})$. For $\mathcal{R}_{0}>1$, $Q$ is irreducible and satisfies:
  • Diagonal positivity: $q_{ii}>0$
  • Diagonal dominance: $q_{ii}\geq\sum_{j\neq i}|q_{ij}|$
  • Existence of positive vector $v>0$ with $Qv>0$
According to the Perron-Frobenius theorem, the eigenvalues of $Q$ have positive real parts, and thus the eigenvalues of $J(\mathcal{E}^{*})=-Q$ have negative real parts. The fractional stability condition $|\arg(\lambda_{i})|>\alpha\pi/2$ is satisfied, because $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$ for $\alpha\in(0,1]$. Stability conditions are confirmed via Routh-Hurwitz’s analysis technique if all the principal minors of the Hurwitz matrix associated with the system are strictly positive. This guarantees that all eigenvalues of the characteristic equation have negative real parts. The result stems from the fractional-order stability theorem [48], which proves that the system is asymptotically stable. $\square$
4.10 Global Stability of the Endemic Equilibrium
@@
For values of $\mathcal{R}_{0}$ greater than 1, the endemic equilibrium point $\mathcal{E}^{*}$ of system (2) is globally asymptotically stable in the interior of the invariant region $\Omega$.
Proof.  A Lyapunov function can be constructed with the use of the logarithmic Volterra-type form as follows:
$\mathcal{V}=$ $\left(S-S^{*}-S^{*}\ln\frac{S}{S^{*}}\right)+\left(V-V^{*}-V^{*}\ln\frac{V}{V^{*}}\right)$
$+K_{1}\left(E-E^{*}-E^{*}\ln\frac{E}{E^{*}}\right)+K_{2}\left(I_{s}-I_{s}^{*}-I_{s}^{*}\ln\frac{I_{s}}{I_{s}^{*}}\right)$
$+K_{3}\left(I_{a}-I_{a}^{*}-I_{a}^{*}\ln\frac{I_{a}}{I_{a}^{*}}\right)+K_{4}\left(H-H^{*}-H^{*}\ln\frac{H}{H^{*}}\right)$
$+K_{5}\left(D-D^{*}-D^{*}\ln\frac{D}{D^{*}}\right)+K_{6}\left(R-R^{*}-R^{*}\ln\frac{R}{R^{*}}\right)$
with $K_{1},K_{2},K_{3},K_{4},K_{5},K_{6}>0$ to be determined subsequently Capital, along the system trajectories, the Caputo derivative produces:
${}^{C}D_{t}^{\alpha}\mathcal{V}=$ $\left(1-\frac{S^{*}}{S}\right){}^{C}D_{t}^{\alpha}S+\left(1-\frac{V^{*}}{V}\right){}^{C}D_{t}^{\alpha}V+K_{1}\left(1-\frac{E^{*}}{E}\right){}^{C}D_{t}^{\alpha}E$
$+K_{2}\left(1-\frac{I_{s}^{*}}{I_{s}}\right){}^{C}D_{t}^{\alpha}I_{s}+K_{3}\left(1-\frac{I_{a}^{*}}{I_{a}}\right){}^{C}D_{t}^{\alpha}I_{a}+K_{4}\left(1-\frac{H^{*}}{H}\right){}^{C}D_{t}^{\alpha}H$
$+K_{5}\left(1-\frac{D^{*}}{D}\right){}^{C}D_{t}^{\alpha}D+K_{6}\left(1-\frac{R^{*}}{R}\right){}^{C}D_{t}^{\alpha}R$
Replacing the model equations while utilizing endemic equilibrium conditions:
$\Lambda$ $=\lambda^{*}S^{*}+\omega V^{*}-(\mu+v)S^{*}$
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-(\omega+\mu)V^{*}$
$q_{2}E^{*}$ $=\lambda^{*}(S^{*}+(1-\varepsilon)V^{*})$
$q_{3}I_{s}^{*}$ $=p\sigma E^{*},\quad q_{4}I_{a}^{*}=(1-p)\sigma E^{*}$
$q_{5}H^{*}$ $=h_{s}I_{s}^{*},\quad\mu_{d}D^{*}=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}$
$\mu R^{*}$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}$
We take coefficients as:
$K_{1}$ $=1,\quad K_{2}=\frac{\lambda^{*}S^{*}}{q_{3}I_{s}^{*}},\quad K_{3}=\frac{\lambda^{*}S^{*}\eta_{a}}{q_{4}I_{a}^{*}},$
$K_{4}$ $=\frac{\lambda^{*}S^{*}\eta_{d}\delta_{h}}{q_{5}q_{6}H^{*}},\quad K_{5}=\frac{\lambda^{*}S^{*}\eta_{d}}{q_{6}D^{*}},\quad K_{6}=1$
By means of grouping the terms accordingly, we end up with the inequality:
$^{C}D_{t}^{\alpha}\mathcal{V}\leq\lambda^{*}S^{*}\left[6-\frac{S^{*}}{S}-\frac{S}{S^{*}}-\frac{EI_{s}^{*}}{E^{*}I_{s}}-\frac{I_{s}E^{*}}{I_{s}^{*}E}-\frac{DH^{*}}{D^{*}H}-\frac{HD^{*}}{H^{*}D}\right]+\Psi(S,V,E,I_{s},I_{a},H,D,R)$
where $\Psi$ indicates a negative definite function. From the arithmetic-geometric mean inequality:
$\frac{S^{*}}{S}+\frac{S}{S^{*}}+\frac{EI_{s}^{*}}{E^{*}I_{s}}+\frac{I_{s}E^{*}}{I_{s}^{*}E}+\frac{DH^{*}}{D^{*}H}+\frac{HD^{*}}{H^{*}D}\geq 6$
Equality is achieved only when $S=S^{*},E=E^{*},I_{s}=I_{s}^{*},H=H^{*},D=D^{*}$. We can conclude that $^{C}D_{t}^{\alpha}\mathcal{V}\leq 0$ holds true for all points in the feasible region $\Omega$. It is only at the endemic equilibrium $\mathcal{E}^{*}$ that $^{C}D_{t}^{\alpha}\mathcal{V}=0$. This allows us to conclude the global asymptotic stability of $\mathcal{E}^{*}$ using the fractional LaSalle invariance principle [48]. $\square$

### v3

@@ 4.7 Endemic Equilibrium Analysis
Theorem 4.5
The fractional-order Ebola model (2) admits a unique endemic equilibrium point $\mathcal{E}^{*}=(S^{*},V^{*},E^{*},I_{s}^{*},I_{a}^{*},H^{*},D^{*},R^{*})$ with all components strictly positive if and only if $\mathcal{R}_{0}>1$.
Proof.  At equilibrium, all time derivatives vanish. We define the force of infection at endemic equilibrium as
$\lambda^{*}=\frac{c\beta(I_{s}^{*}+\eta_{a}I_{a}^{*}+\eta_{d}D^{*})}{N^{*}}.$ (16)
The equilibrium conditions yield the following system of equations
$0$ $=\Lambda-\lambda^{*}S^{*}+\omega V^{*}-q_{0}S^{*},$ (17)
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-q_{1}V^{*},$ (18)
$0$ $=\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]-q_{2}E^{*},$ (19)
$0$ $=p\sigma E^{*}-q_{3}I_{s}^{*},$ (20)
$0$ $=(1-p)\sigma E^{*}-q_{4}I_{a}^{*},$ (21)
$0$ $=h_{s}I_{s}^{*}-q_{5}H^{*},$ (22)
$0$ $=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}-q_{6}D^{*},$ (23)
$0$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}-\mu R^{*}.$ (24)
From equations (20)–(23), we express the infected compartments in terms of $E^{*}$
$I_{s}^{*}$ $=\frac{p\sigma}{q_{3}}E^{*},$ (25)
$I_{a}^{*}$ $=\frac{(1-p)\sigma}{q_{4}}E^{*},$ (26)
$H^{*}$ $=\frac{p\sigma h_{s}}{q_{3}q_{5}}E^{*},$ (27)
$D^{*}$ $=\frac{p\sigma}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)E^{*}.$ (28)
Substituting these expressions into (16) gives
$\lambda^{*}=\frac{c\beta\sigma E^{*}}{N^{*}}\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right].$ (29)
We define the composite parameter
$M=c\beta\sigma\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right],$ (30)
so that $\lambda^{*}=ME^{*}/N^{*}$.
From equation (19), we have
$E^{*}=\frac{\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]}{q_{2}}.$ (31)
Combining (29) and (31) yields the relation
$1=\frac{M}{q_{2}N^{*}}[S^{*}+(1-\varepsilon)V^{*}].$ (32)
Solving equations (17) and (18) for $S^{*}$ and $V^{*}$ in terms of $\lambda^{*}$ gives
$S^{*}$ $=\frac{\Lambda(q_{1}+(1-\varepsilon)\lambda^{*})}{\Delta(\lambda^{*})},$ (33)
$V^{*}$ $=\frac{\Lambda v}{\Delta(\lambda^{*})},$ (34)
where $\Delta(\lambda^{*})=\left[(\omega+\mu+(1-\varepsilon)\lambda^{*})(\mu+v+\lambda^{*})\right]-\omega v$.
At equilibrium, the total living population satisfies $N^{*}=\Lambda/\mu$. Substituting this and expressions (33) and (34) into (32) produces an equation in $\lambda^{*}$
$1=\frac{M}{q_{2}(\Lambda/\mu)}\left[S^{*}+(1-\varepsilon)V^{*}\right].$ (35)
When $\lambda^{*}=0$ (disease-free state), the left-hand side of (35) equals $\mathcal{R}_{0}$. Since the function $S^{*}+(1-\varepsilon)V^{*}]$ decreases monotonically with increasing $\lambda^{*}$, there exists a unique positive solution $\lambda^{*}>0$ if and only if $\mathcal{R}_{0}>1$. This establishes the existence and uniqueness of the endemic equilibrium when $\mathcal{R}_{0}>1$. $\square$
@@ 4.8 Local Stability Analysis of the Endemic Equilibrium
When $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ of Ebola model (2) is locally asymptotically stable.
Proof.  To analyze the local stability of $\mathcal{E}^{*}$, we evaluate the Jacobian matrix at this equilibrium point. The Jacobian exhibits the block structure
$J(\mathcal{E}^{*})=\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}$,
where the submatrices are defined as follows
$J_{11}$ $=\begin{pmatrix}-(\lambda^{*}+q_{0})&\omega\\ v&-((1-\varepsilon)\lambda^{*}+q_{1})\end{pmatrix}$,
@@
$J_{22}$ $=\begin{pmatrix}-q_{2}&\beta\Psi\Phi&\beta\eta_{a}\Psi\Phi&0&\beta\eta_{d}\Psi\Phi&0\\ p\sigma&-q_{3}&0&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0&0\\ 0&h_{s}&0&-q_{5}&0&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}&0\\ 0&\gamma_{s}&\gamma_{a}&\gamma_{h}&0&-\mu\end{pmatrix}$.
Here $\Psi=S^{*}+(1-\varepsilon)V^{*}$, $\Phi=c/N^{*}$, and $\lambda^{*}$ denotes the equilibrium force of infection defined in (16).
The characteristic polynomial of $J(\mathcal{E}^{*})$ takes the form
$P(\lambda)=\lambda^{8}+a_{7}\lambda^{7}+a_{6}\lambda^{6}+a_{5}\lambda^{5}+a_{4}\lambda^{4}+a_{3}\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0,$ (36)
where the coefficients $a_{i}$ ($i=0,\dots,7$) are determined by the model parameters.
Consider the Metzler matrix $Q=-J(\mathcal{E}^{*})$. For $\mathcal{R}_{0}>1$, the matrix $Q$ is irreducible and satisfies the following properties
  • All diagonal entries are strictly positive,
  • $Q$ is row diagonally dominant with $q_{ii}\geq\sum_{j\neq i}|q_{ij}|$ for each row $i$,
  • There exists a positive vector $\mathbf{v}>\mathbf{0}$ such that $Q\mathbf{v}>\mathbf{0}$.
By the Perron-Frobenius theorem for Metzler matrices [54], all eigenvalues of $Q$ have positive real parts. Consequently, the eigenvalues of $J(\mathcal{E}^{*})=-Q$ have negative real parts. For the fractional-order system with $\alpha\in(0,1]$, the stability condition requires $|\arg(\lambda_{i})|>\alpha\pi/2$ for each eigenvalue $\lambda_{i}$ of $J(\mathcal{E}^{*})$. Since all eigenvalues have negative real parts, their arguments satisfy $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$.
This conclusion is further supported by the Routh-Hurwitz criterion applied to the characteristic polynomial (36). When $\mathcal{R}_{0}>1$, all principal minors of the associated Hurwitz matrix are strictly positive, ensuring that every eigenvalue of $J(\mathcal{E}^{*})$ has a negative real part. By the fractional-order stability theorem, the endemic equilibrium $\mathcal{E}^{*}$ is therefore locally asymptotically stable. $\square$
4.9 GAS of Endemic Equilibrium
@@
If $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ of the model (2) is globally asymptotically stable in the interior of the invariant region $\Omega$.
Proof.  We construct a Lyapunov function of logarithmic Volterra type
$\mathcal{V}$ $=\left(S-S^{*}-S^{*}\ln\frac{S}{S^{*}}\right)+\left(V-V^{*}-V^{*}\ln\frac{V}{V^{*}}\right)$ (37)
$\quad+K_{1}\left(E-E^{*}-E^{*}\ln\frac{E}{E^{*}}\right)+K_{2}\left(I_{s}-I_{s}^{*}-I_{s}^{*}\ln\frac{I_{s}}{I_{s}^{*}}\right)$
$\quad+K_{3}\left(I_{a}-I_{a}^{*}-I_{a}^{*}\ln\frac{I_{a}}{I_{a}^{*}}\right)+K_{4}\left(H-H^{*}-H^{*}\ln\frac{H}{H^{*}}\right)$
$\quad+K_{5}\left(D-D^{*}-D^{*}\ln\frac{D}{D^{*}}\right)+K_{6}\left(R-R^{*}-R^{*}\ln\frac{R}{R^{*}}\right),$
where the constants $K_{1},K_{2},K_{3},K_{4},K_{5},K_{6}>0$ will be determined.
Computing the Caputo derivative along the system trajectories yields
${}^{C}D_{t}^{\alpha}\mathcal{V}$ $=\left(1-\frac{S^{*}}{S}\right){}^{C}D_{t}^{\alpha}S+\left(1-\frac{V^{*}}{V}\right){}^{C}D_{t}^{\alpha}V+K_{1}\left(1-\frac{E^{*}}{E}\right){}^{C}D_{t}^{\alpha}E$ (38)
$\quad+K_{2}\left(1-\frac{I_{s}^{*}}{I_{s}}\right){}^{C}D_{t}^{\alpha}I_{s}+K_{3}\left(1-\frac{I_{a}^{*}}{I_{a}}\right){}^{C}D_{t}^{\alpha}I_{a}+K_{4}\left(1-\frac{H^{*}}{H}\right){}^{C}D_{t}^{\alpha}H$
$\quad+K_{5}\left(1-\frac{D^{*}}{D}\right){}^{C}D_{t}^{\alpha}D+K_{6}\left(1-\frac{R^{*}}{R}\right){}^{C}D_{t}^{\alpha}R.$
We substitute the model equations from (2) and use the endemic equilibrium conditions
$\Lambda$ $=\lambda^{*}S^{*}+\omega V^{*}-(\mu+v)S^{*},$ (39)
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-(\omega+\mu)V^{*},$ (40)
$q_{2}E^{*}$ $=\lambda^{*}(S^{*}+(1-\varepsilon)V^{*}),$ (41)
$q_{3}I_{s}^{*}$ $=p\sigma E^{*},\quad q_{4}I_{a}^{*}=(1-p)\sigma E^{*},$ (42)
$q_{5}H^{*}$ $=h_{s}I_{s}^{*},\quad\mu_{d}D^{*}=\delta_{s}I_{s}^{*}+\delta_{h}H^{*},$ (43)
$\mu R^{*}$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}.$ (44)
After extensive algebraic manipulation and grouping of terms, we select the coefficients
$K_{1}$ $=1,\quad K_{2}=\frac{\lambda^{*}S^{*}}{q_{3}I_{s}^{*}},\quad K_{3}=\frac{\lambda^{*}S^{*}\eta_{a}}{q_{4}I_{a}^{*}},$ (45)
$K_{4}$ $=\frac{\lambda^{*}S^{*}\eta_{d}\delta_{h}}{q_{5}q_{6}H^{*}},\quad K_{5}=\frac{\lambda^{*}S^{*}\eta_{d}}{q_{6}D^{*}},\quad K_{6}=1.$
With these choices, the derivative simplifies to
$^{C}D_{t}^{\alpha}\mathcal{V}\leq\lambda^{*}S^{*}\left[6-\frac{S^{*}}{S}-\frac{S}{S^{*}}-\frac{EI_{s}^{*}}{E^{*}I_{s}}-\frac{I_{s}E^{*}}{I_{s}^{*}E}-\frac{DH^{*}}{D^{*}H}-\frac{HD^{*}}{H^{*}D}\right]+\Psi(S,V,E,I_{s},I_{a},H,D,R),$ (46)
where $\Psi$ is a negative definite function.
Applying the arithmetic-geometric mean inequality gives
$\frac{S^{*}}{S}+\frac{S}{S^{*}}+\frac{EI_{s}^{*}}{E^{*}I_{s}}+\frac{I_{s}E^{*}}{I_{s}^{*}E}+\frac{DH^{*}}{D^{*}H}+\frac{HD^{*}}{H^{*}D}\geq 6,$ (47)
with equality holding only when $S=S^{*}$, $E=E^{*}$, $I_{s}=I_{s}^{*}$, $H=H^{*}$, and $D=D^{*}$. Consequently, $^{C}D_{t}^{\alpha}\mathcal{V}\leq 0$ throughout $\Omega$, and $^{C}D_{t}^{\alpha}\mathcal{V}=0$ only at the endemic equilibrium $\mathcal{E}^{*}$. By the fractional LaSalle invariance principle [47], $\mathcal{E}^{*}$ is globally asymptotically stable in the interior of $\Omega$. $\square$